OpenStudy (anonymous):

Select one of the factors of x3 + 4x2 + 9x + 36.

myininaya (myininaya):

Try factoring by grouping!

myininaya (myininaya):

What do the first two terms have in common?

OpenStudy (anonymous):

x3 and 9x?

myininaya (myininaya):

\[x^3 \text{ and } 4x^2 \text{ are the first two terms}\]

myininaya (myininaya):

What do they have in common?

myininaya (myininaya):

\[x^3=x \cdot x^2\] \[4x^2=4 \cdot x^2\] Maybe this will make it more obvious?

OpenStudy (anonymous):

the variable x

myininaya (myininaya):

They both have \[x^2\] in common

myininaya (myininaya):

So we can write \[x^2(x+4)+9x+36\] Now what do the last two terms have in common?

myininaya (myininaya):

In other words what I'm asking you to find is this thing called the gcf (Greatest common factor) of 9x and 36?

OpenStudy (anonymous):

the gcf is 3

myininaya (myininaya):

what about 9?

myininaya (myininaya):

\[x^2(x+4)+9(x+4)\] Now you have two terms and they both have what in common?